# Efficiency vs. Efficacy: Assessing the Compression Ratio-Dice Score Relationship through a Simple Benchmarking Framework for Cerebrovascular 3D Segmentation

### v2

@@ 1 Introduction
In light of this, our study makes several key contributions to the research community. While prior work has explored compression and segmentation separately, and some have investigated the impact of compression on generic image quality, our research directly and quantitatively assesses its effects on a specific, critical downstream task: automated cerebrovascular segmentation. We are aware of a diverse set of techniques for 3D medical volume compression including classical techniques such as ZFP and more advanced techniques such as autoencoder-based compressors [1, 2, 3] and Implicit Neural Representations (INRs) [4, 5, 6]. However, providing a comprehensive, head-to-head comparison of the performance of these families of techniques is beyond the scope of this research. We strictly focus on the ZFP compression.
@@ 2 Related Work
In [7], this research focuses on enhancing the JP3D compression standard—the 3D extension of JPEG 2000, which relies on the Discrete Wavelet Transform (DWT). The authors introduce novel methods, including hybrid adaptive transforms and histogram packing, to overcome the limitations of fixed-basis wavelets and improve compression ratios (CRs) for various volumetric modalities (CT, MRI, Ultrasound). They demonstrate CR improvements of up to 6.5% at an acceptable cost, but their evaluation primarily centers on file size reduction and maximizing compression efficiency rather than assessing the impact of the lossy process on critical downstream diagnostic tasks like segmentation.
@@
The authors of [8] proposed an extension of the 3D Wavelet Difference Reduction (3D-WDR) algorithm, specifically tailored for volumetric data by exploiting spatiotemporal coherence between adjacent slices. The method utilizes metrics like Mean Co-located Pixel Difference (MCPD) to optimally group and compress similar slices, achieving high compression rates (down to one bit per voxel). The performance is rigorously evaluated using traditional image quality metrics, such as Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM), with the explicit goal of maintaining high visual quality for human diagnosis, making it a representative example of image-centric compression research.
@@
The research presented in [9] attempted to modernize transform-based coding by introducing a versatile, end-to-end framework called aiWave, which employs a trained 3D affine wavelet-like transform based on the lifting scheme. The goal is to overcome the performance ceilings imposed by the manually-designed wavelet bases used in standards like JP3D. While the use of a learned component blurs the line between traditional and deep learning approaches, the core mechanism remains transform-based, and the evaluation is primarily focused on rate-distortion performance, showcasing the frontier of general-purpose volumetric coding before task-specific metrics are introduced.
@@
The research proposed in [10] directly tackles the incompatibility between compression optimized for the Human Visual System (HVS) and the requirements of Deep Neural Networks (DNNs) used for cloud-based segmentation. The authors demonstrate that traditional compression (like JPEG 2000) that looks visually acceptable to a radiologist can critically degrade the accuracy of an automated segmentation task. They propose a novel framework tailored for machine vision that uses task-specific feature extraction to automatically retain only the image features crucial for accurate segmentation, ultimately achieving higher segmentation accuracy (Dice score) at equivalent or better compression rates than existing general-purpose methods.
@@
This study presented in [11] investigated the robustness of standard segmentation networks, such as the U-Net architecture, against artifacts introduced by aggressive lossy compression using the JPEG 2000 standard on multiple 3D medical datasets. Contrary to the assumption that any lossy compression harms diagnostic performance, the authors empirically demonstrate that DNN models trained on compressed images remain remarkably robust, showing no significant drop in segmentation accuracy (Dice score) even at high compression rates (up to 20 times). This suggests that for cloud-based training and inference, substantial data reduction is possible using existing compression methods without compromising deep learning model performance.
@@
The authors of [12] introduced an advanced end-to-end Task-Driven Semantic-Aware Image Compression (TDSIC) framework specifically designed for Internet of Things/Everything (IoT/IoE) applications where downstream analysis tasks are prioritized over mere human visual perception. TDSIC achieves this by integrating several novel mechanisms: using importance maps as prior knowledge to guide bit allocation to semantically crucial regions, refining the perceptual loss function to prioritize task-relevant features, and incorporating a feature enhancement network (FEN) to effectively recover latent semantic information lost during compression. By utilizing feedback from discriminative networks to calibrate the perceptual model and successfully recovering lost information, TDSIC demonstrates superior performance, achieving up to a 12% improvement in Top-1 classification accuracy and significant gains in object detection, while simultaneously preserving acceptable visual quality for human viewers.
@@ 3.3 Segmentation and Evaluation
For the automated segmentation, we used a flavor of the MambaVesselNet architecture proposed in [13]. This is a hybrid network which successfully integrates the strengths of Convolutional Neural Networks (CNNs)—excellent for capturing local, low-level semantic features—with the Mamba blocks—which efficiently model long-range dependencies across the entire 3D volume with linear complexity. This combination is particularly effective for segmenting complex, elongated structures like cerebrovascular networks where maintaining topological connectivity and global context is essential, overcoming the high computational cost limitations typically faced by transformer-based models in 3D medical imaging. This architecture was adopted in this research for being a state-of-the-art model for volumetric medical image segmentation.
The model was trained on the original, uncompressed volumes with their corresponding ground-truth segmentation masks. The trained model was then used to infer segmentation masks on both the uncompressed volumes and the compressed/decompressed versions. To quantitatively assess the impact of compression, we used widely-accepted metrics to compare the inferred segmentation masks to the ground truth. These include the following: 1) the Dice Similarity Coefficient (DSC) 2) and the Intersection over Union (IoU).
The Dice Similarity Coefficient (Dice) and the Jaccard Index (IoU) are the two primary metrics for evaluating segmentation quality, both measuring the overlap between the predicted mask and the ground truth. While both metrics are fundamentally and mathem atically related, they exhibit different sensitivities: the Dice score is generally more forgiving of false positives and is commonly reported, whereas the IoU metric penalizes large differences more heavily, providing a stricter measure of overlap, particularly in small or thin structures. Reporting both Dice and IoU is crucial as it offers a more comprehensive assessment of segmentation performance, confirming both the general accuracy and the precision required for clinically relevant boundary delineation.
@@ 4 Experimental Results
| Rate (bits/voxel) | Overall Mean DICE | Overall Mean IoU | Overall Mean Sensitivity | Overall Mean Volume Similarity |
@@
| Abs. Error Tolerance | Overall Mean DICE | Overall Mean IoU | Overall Mean Sensitivity | Overall Mean Volume Similarity |

### v3

@@ 1 Introduction
In light of this, our study makes several key contributions to the research community. While prior work has explored compression and segmentation separately, and some have investigated the impact of compression on generic image quality, our research directly and quantitatively assesses its effects on a specific, critical downstream task: automated cerebrovascular segmentation. We are aware of a diverse set of techniques for 3D medical volume compression including classical techniques such as ZFP and more advanced techniques such as autoencoder-based compressors [varma2025medvae, li20243d, thadikemalla20243d] and Implicit Neural Representations (INRs) [sheibanifard2025end, li2025compact, vzuurkova2025accelerated]. However, providing a comprehensive, head-to-head comparison of the performance of these families of techniques is beyond the scope of this research. We strictly focus on the ZFP compression.
@@ 2 Related Work
In [starosolski2020employing], this research focuses on enhancing the JP3D compression standard—the 3D extension of JPEG 2000, which relies on the Discrete Wavelet Transform (DWT). The authors introduce novel methods, including hybrid adaptive transforms and histogram packing, to overcome the limitations of fixed-basis wavelets and improve compression ratios (CRs) for various volumetric modalities (CT, MRI, Ultrasound). They demonstrate CR improvements of up to 6.5% at an acceptable cost, but their evaluation primarily centers on file size reduction and maximizing compression efficiency rather than assessing the impact of the lossy process on critical downstream diagnostic tasks like segmentation.
@@
The authors of [zerva2020improving] proposed an extension of the 3D Wavelet Difference Reduction (3D-WDR) algorithm, specifically tailored for volumetric data by exploiting spatiotemporal coherence between adjacent slices. The method utilizes metrics like Mean Co-located Pixel Difference (MCPD) to optimally group and compress similar slices, achieving high compression rates (down to one bit per voxel). The performance is rigorously evaluated using traditional image quality metrics, such as Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM), with the explicit goal of maintaining high visual quality for human diagnosis, making it a representative example of image-centric compression research.
@@
The research presented in [xue2022aiwave] attempted to modernize transform-based coding by introducing a versatile, end-to-end framework called aiWave, which employs a trained 3D affine wavelet-like transform based on the lifting scheme. The goal is to overcome the performance ceilings imposed by the manually-designed wavelet bases used in standards like JP3D. While the use of a learned component blurs the line between traditional and deep learning approaches, the core mechanism remains transform-based, and the evaluation is primarily focused on rate-distortion performance, showcasing the frontier of general-purpose volumetric coding before task-specific metrics are introduced.
@@
The research proposed in [liu2019machine] directly tackles the incompatibility between compression optimized for the Human Visual System (HVS) and the requirements of Deep Neural Networks (DNNs) used for cloud-based segmentation. The authors demonstrate that traditional compression (like JPEG 2000) that looks visually acceptable to a radiologist can critically degrade the accuracy of an automated segmentation task. They propose a novel framework tailored for machine vision that uses task-specific feature extraction to automatically retain only the image features crucial for accurate segmentation, ultimately achieving higher segmentation accuracy (Dice score) at equivalent or better compression rates than existing general-purpose methods.
@@
This study presented in [kurmukov2024effect] investigated the robustness of standard segmentation networks, such as the U-Net architecture, against artifacts introduced by aggressive lossy compression using the JPEG 2000 standard on multiple 3D medical datasets. Contrary to the assumption that any lossy compression harms diagnostic performance, the authors empirically demonstrate that DNN models trained on compressed images remain remarkably robust, showing no significant drop in segmentation accuracy (Dice score) even at high compression rates (up to 20 times). This suggests that for cloud-based training and inference, substantial data reduction is possible using existing compression methods without compromising deep learning model performance.
@@
The authors of [zhou2024task] introduced an advanced end-to-end Task-Driven Semantic-Aware Image Compression (TDSIC) framework specifically designed for Internet of Things/Everything (IoT/IoE) applications where downstream analysis tasks are prioritized over mere human visual perception. TDSIC achieves this by integrating several novel mechanisms: using importance maps as prior knowledge to guide bit allocation to semantically crucial regions, refining the perceptual loss function to prioritize task-relevant features, and incorporating a feature enhancement network (FEN) to effectively recover latent semantic information lost during compression. By utilizing feedback from discriminative networks to calibrate the perceptual model and successfully recovering lost information, TDSIC demonstrates superior performance, achieving up to a 12% improvement in Top-1 classification accuracy and significant gains in object detection, while simultaneously preserving acceptable visual quality for human viewers.
@@ 3.3 Segmentation and Evaluation
For volumetric cerebrovascular segmentation, we adopted two state-of-the-art Mamba-based architectures: MambaVesselNet [chen2024mambavesselnet] and U-Mamba [ma2024u]. MambaVesselNet is a hybrid framework that combines the strong local feature extraction capabilities of Convolutional Neural Networks (CNNs) with the efficient long-range dependency modeling of Mamba blocks, enabling accurate segmentation of complex, elongated vascular structures while avoiding the high computational costs typically associated with 3D Transformer models. Similarly, U-Mamba extends this design philosophy by integrating CNNs with State Space Sequence Models (SSMs) to overcome CNNs’ inherent locality and Transformers’ quadratic attention complexity, achieving linear-complexity global context modeling suitable for biomedical images. In our study, the MambaVesselNet variant was trained on full-resolution volumes and corresponding ground-truth masks, and subsequently used to infer segmentations on both original and compressed/decompressed volumes. Segmentation performance was quantitatively assessed using two widely accepted metrics—the Dice Similarity Coefficient (DSC) and the Intersection over Union (IoU). While mathematically related, the Dice score is more tolerant to false positives, whereas IoU penalizes discrepancies more strictly, especially in thin or small structures. Reporting both metrics provides a comprehensive and clinically relevant evaluation of segmentation accuracy and boundary precision.
@@ 4 Experimental Results
| Model | Rate (bits/voxel) | Overall | Mean | Overall | Mean |
| DICE | IoU | Sensitivity | Volume Similarity | ||
| U-Mamba | Original CT | ||||
| 2 | |||||
| 4 | |||||
| 8 | |||||
| 16 | |||||
| MambaVesselNet++ | 2 | ||||
| 4 | |||||
| 8 | |||||
| 16 |
@@
| Model | Abs. Error Tolerance | Overall DICE | Mean IoU | Overall Sensitivity | Mean Volume Similarity |
| MambaVesselNet | 500 | ||||
| 1000 | |||||
| 1500 | |||||
| U-Mamba | Original CT | ||||
| 500 | |||||
| 1000 | |||||
| 1500 |